OpenStudy (anonymous):

What is the value of x? A. 5.7 B. 8.5 C. 9.0 D. 17.0

OpenStudy (zzr0ck3r):

um we need more information

OpenStudy (zzr0ck3r):

this is sort of like asking what color is the car? red blue green ??????????????????????// what car?

OpenStudy (anonymous):

solve \[4x-16=2x-8\] for \(x\)

OpenStudy (anonymous):

thx btw satellite i'm just trying to ask a question PLEASE DON'T SUSPEND ME

OpenStudy (anonymous):

Thx

OpenStudy (anonymous):

wait that's not it

OpenStudy (anonymous):

4x-26=2x-8

OpenStudy (anonymous):

That is your equation

OpenStudy (anonymous):

Sorry, the question said 4x-16=2x-8, so that's how I got 4.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

4x-26=2x-8 -2x -2x 2x-26=-8 +26 +26 2x=18

OpenStudy (anonymous):

@KingGeorge plz don't suspend me this time i'm just trying to ask a question I REALLY REALLY need answered

OpenStudy (anonymous):

divide 18 by 2

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

and you will get your answer

OpenStudy (anonymous):

c is the answer!

OpenStudy (anonymous):

@arilove1d I still have a few more questions though

OpenStudy (anonymous):

Correct ! 9.0 is the answer !:)

OpenStudy (anonymous):

Ok ill help you with as much as i can

OpenStudy (anonymous):

__ What is the length of BC approximated to the nearest tenth? A. 11 B. 13 C. 14 D. 16

OpenStudy (anonymous):

Please help

OpenStudy (anonymous):

@arilove1d

hartnn (hartnn):

the " | " shown at those sides imply that the two sides are equal in length so, equating their lengths, we get, 6x-50 = x+5 solving this for x, 6x - x = 5 + 50 5x = 55 x=11 :)

OpenStudy (anonymous):

A